OpenStudy (anonymous):

Solve: x+y=5 2x-3y=0

OpenStudy (anonymous):

x=5-y so sub this into the second equation. 2(5-y) - 3y = 0 10-2y - 3y = 0 10 - 5y = 0 10 = 5y 2 = y

OpenStudy (anonymous):

what is x?

OpenStudy (anonymous):

sorry, just put y = 2 into the first equation.. so.. x = 5 - (2) = 3

OpenStudy (anonymous):

You know that y=2 right so... You plug it into the equation x=5-y Which will become x=5-2